five pens of the same type cost the same as two notebooks of the same type.If one pen and 2 notebooks cost$4.20 how much does one one pen cost?
Answer by oberobic(2304)

5p = 2n, where p = cost of pens and n = cost of notebooks
so,
n = 2.5p
.
p + 2n = 4.20, which is given
.
substituting
p + 2(2.5p) = 4.20
p + 5p = 4.20
6p = 4.20
p = .70 = 70 cents = cost of one pen
.
We also can calculate the cost of one notebook.
5p = 2n
5(.70) = 2n
3.50 = 2n
n = 1.75
.
Done